Word Problem help!!
A student makes a $9.75 purchase at the book store with a $20 bill. The store has no bills and gives the change in quarters and fifty-cent pieces. There are 30 coins in all. How many of each kind are there? How many quarter are there in the change____?

Answer by mananth(16946)

purchase = $9.75
amount given = $20
balance to return = $10.25
...
let quarters be x
fifty cent will be the balance = 30-x
...
25x+50(30-x)=1025 cents
25x+1500-50x=1025
-25x+1500 =1025
-1500
-25x=1025-1500
-25x=-475
/-25
x=19 quarters
50 cent pieces will be 30-19=11
...
Check
11*50+19*25= 10.25 = $10.25
...
m.ananth@hotmail.ca